OpenStudy (anonymous):

f(x) = sin(sinx^3) f'(x) = ?

OpenStudy (thomas5267):

Have you learn chain rule yet?

OpenStudy (anonymous):

how did you know it was chain rule and not the product rule ?

OpenStudy (abb0t):

chain rule!

OpenStudy (zzr0ck3r):

composition of functions uses chain rule multiplication of functions uses the product rule

OpenStudy (thomas5267):

Because you have a composition of function instead of multiplication. If you have \(f(x)g(x)\), then use product rule. If you have \(f(g(x))\), use chain rule instead.

OpenStudy (anonymous):

cos(sinx^3)(3cosx^2) ?

OpenStudy (anonymous):

\[\begin{split} d(\sin(\sin(x^3))) &= \cos(\sin(x^3))d(\sin(x^2)) \\ &= \cos(\sin(x^3))\cos(x^3)d(x^3) \\ &= \cos(\sin(x^3))\cos(x^3)3x^2dx \end{split} \]

OpenStudy (anonymous):

oh pellet.

OpenStudy (anonymous):

thanks <3

OpenStudy (thomas5267):

It is interesting how sh!t becomes "pellet".

OpenStudy (anonymous):

Lol yea